A sample of purchases at the local convenience store has resulted in the following sample data, where x = the number of items purchased per customer and f = the number of customers.
x 1 2 3 4 5

f 7 17 8 8 7
What does the 17 stand for in the above table?
Find the midrange of items purchased.
How many items were purchased by the customers in this sample?

Answer by stanbon(75887)

A sample of purchases at the local convenience store has resulted in the following sample data, where x = the number of items purchased per customer and f = the number of customers.
x 1 2 3 4 5
f 7 17 8 8 7
What does the 17 stand for in the above table?
17 of the customers purchased 2 items.
-------------------------------------------
Find the midrange of items purchased.
Lower range: 7*1+17*2+8*3 = 65
Upper range: 7*5+8*4*8*3 = 91
Midrange of items purchased is 3
--------------------------------------
How many items were purchased by the customers in this sample?
Total = 7*1 + 17*2 + 8*3 etc.
-----
Cheers,
Stan H.